Discuss in terms of the physical properties of a capacitor whether more or less energy is stored in a capacitor with a dielectric constant greater than 1 as compared to an air filled capacitor?
(Please be elaborative)
Answer #1

The energy stored in an air filled capacitor (with dielectrec constant \simeq 1) is similar with a theorical void capacitor:

U_o = \frac{1}{2} \frac{Q^2}{C_o}

For a dielectric with constant grater than 1

K > 1 \Rightarrow K=\frac{C}{C_o}>1 \Rightarrow C>C_o

that means that in a capacitor filled with dielectric with constant greater than 1 the capacitance is greater than in an filled with air capacitor. The energy in the capacitor with a dielectric will be

U = \frac{1}{2}\frac{Q^2}{C}

Since C >C_o the energy will be U_o > U. In other words, the energy in the caparcitor filled with a dielectric with constant greaster than 1 (U) is less than the energy in an air filled capacitor (U_o) with constant almost 1.

The energy is less because the polarization in the dielectric.